On 1/1/2016, Wilson buys 25% of Taylormade for $500,000 and accounts for the

investment using the equity method. The carrying value of net assets at date of purchase

totaled $1,800,000. Fair value and the carrying values of the acquired assets were same

except for a factory with a 10 year useful life, which exceeded its carrying value by

$150,000 and inventory which exceeded its carrying value by $50,000. The inventory

was sold during 2016. (That is the fair value of Taylormade’s net assets were

$2,000,000). During 2016, Taylormade reports net income of $200,000 and paid $40,000

cash dividend. What amount should Wilson report as the carrying amount for its

investment in Taylormade at December 31

st

, 2016.

a.

$595,000

b.

$532,000

c.

$493,750

d.

$523,750
Answer #1

Solution:

Adjusted net income for Taylormade considering fair value of asset considered at the time of acquisition = $200,000 - Inventory sold having fair value exceeded by $50,000 - Depreciation on factory for difference in fair value and carrying value

= $200,000 - $50,000 - ($150,000/10) = $135,000

Share of income in taylor made to be recognized by Wilson = $135,000* 25% = $33,750

Share of dividend received by wilson = $40,000*25% = $10,000

Wilson report as the carrying amount for its investment in Taylormade at December 31 = Cost of investment + share of income -Share in dividend

= $500,000 + $33,750 - $10,000

= $523,750

Hence option d is correct.